# Evaluating Open Language Models Across Task Types, Application Domains, and Reasoning Types: An In-Depth Experimental Analysis

### v1

@@ Appendix B Aspect-level Analyses
B.1 Performance of LMs on all Aspects
In this appendix, we report results of all 13 LMs (5 pre-trained, 5 IT and 3 SOTA models that we compared our work to) on all entities of all three aspects present in the test set of the dataset. It includes the ones not covered in Section 2.4, but were available in the test-set of Super-Natural Instructions (Wang et al., 2022), with English as the input and output languages. Note, we also provide the results SOTA models for comparisons. Table 5 reports the results for all task types, Table 7 reports the results on all application domains and Table 6 for all reasoning types. Note, we abbreviate the model names at some places in the columns of these tables. The abbreviations and full model names can be found in table 4.
There can be some tasks which can be classified into two aspects, like title generation for News articles will belong to title generation task type and News domain. In all other analyses, each domain has been considered independent, which is not always the case. However, in the dataset, there are many pairwise aspects that do not contain any tasks, and for most of the ones that were present, Mistral-7B-I was the best model. Thus, we are not reporting the tabulated results for those considering the sparsity and repetitiveness of such a dense table. This is also discussed in Section 6.
@@ G-7B-I Gemma-7B-I
Task Type Pre-trained Models Instruction-tuned Models SOTA Models
G-2B M-7B G-7B L-3-8B F-2-11B G-2B-I P-3-I M-7B-I G-7B-I L-3-8B-I GPT-3.5-T GPT-4o DS-2
Answerability Classification 88.80 60.56 88.76 87.50 88.29 91.87 61.01 97.39 86.21 85.98 98.53 98.87 91.78
Cause Effect Classification 86.65 72.47 82.16 82.30 88.00 88.03 68.62 94.40 85.56 85.73 95.58 96.59 91.58
Coreference Resolution 85.26 65.72 75.20 78.76 84.70 88.00 68.45 95.85 84.08 83.27 96.60 97.99 93.12
Data to Text 87.19 73.06 82.01 86.71 88.53 87.72 84.25 90.83 86.14 85.87 91.55 91.94 91.73
Dialogue Act Recognition 86.16 66.20 84.61 83.70 86.60 88.04 63.07 97.04 84.28 83.66 96.73 97.70 91.00
Grammar Error Correction 93.52 90.02 86.47 89.76 87.45 93.72 73.31 95.29 89.25 87.84 95.69 95.79 94.51
Keyword Tagging 84.23 72.79 75.17 83.97 86.03 86.01 71.60 88.52 84.42 83.64 92.08 91.43 89.57
Overlap Extraction 85.03 85.57 84.03 85.27 86.66 86.34 90.08 90.69 85.40 83.83 90.93 92.16 90.04
Question Rewriting 90.05 59.48 87.81 87.21 89.41 91.05 73.09 93.17 87.80 86.85 93.96 94.37 93.46
Textual Entailment 87.91 72.20 82.52 83.96 86.61 88.43 52.79 97.07 84.76 84.41 96.64 99.04 89.44
Title Generation 84.81 57.64 75.57 77.12 86.61 86.67 72.59 89.43 85.44 84.45 88.86 90.01 88.95
Word Analogy 82.67 80.14 83.35 83.22 81.15 88.35 81.26 89.00 85.17 81.53 94.85 97.37 88.44
@@
Reasoning Type Pre-trained Models Instruction-tuned Models SOTA Models
G-2B M-7B G-7B L-3-8B F-2-11B G-2B-I P-3-128k-I M-7B-I G-7B-I L-3-8B-I GPT-3.5-T GPT-4o DS-2
Abductive 87.58 51.75 86.86 86.08 89.67 88.40 82.56 91.36 87.54 87.43 92.11 92.11 91.08
Analogical 83.02 77.24 83.71 83.76 81.73 88.65 82.21 89.15 85.32 81.92 94.61 96.95 88.91
Causal 87.00 78.89 81.19 79.85 86.90 85.18 60.11 93.35 84.78 84.65 94.21 95.88 90.16
Commonsense 86.91 67.08 83.75 82.26 85.61 88.96 65.28 94.31 85.07 83.89 95.98 97.82 91.35
Cross-document 94.98 63.21 88.45 80.34 92.52 92.94 91.94 95.33 84.15 84.73 95.25 95.53 89.10
Deductive 91.14 87.34 78.63 85.61 88.22 89.78 60.13 92.70 83.73 84.49 92.04 94.75 88.18
Discrete 94.98 63.21 88.45 80.34 92.52 92.94 91.94 95.33 84.15 84.73 95.25 95.53 89.10
Logical 83.58 77.17 77.46 80.16 79.78 86.13 26.97 96.19 81.19 80.50 96.88 99.12 88.09
Multihop 94.98 68.29 86.04 79.84 92.07 92.94 90.71 94.11 83.22 83.82 93.75 94.08 89.01
Numerical 94.98 63.21 88.45 80.34 92.52 92.94 91.94 95.33 84.15 84.73 95.25 95.53 89.10
Quantitative 88.67 77.48 83.60 78.10 86.59 84.59 67.94 93.96 84.80 84.46 93.05 95.19 89.15
Reasoning on Actions 85.03 82.10 82.07 86.80 89.73 87.18 74.77 92.31 86.03 85.12 91.65 92.78 92.48
Reasoning on Objects 90.79 92.76 86.00 91.66 92.66 91.81 88.87 93.08 87.16 87.68 93.61 93.64 93.12
Social Interactions 84.63 62.88 83.36 86.33 86.14 87.61 64.95 93.91 84.52 84.33 96.35 96.55 92.32
Relational 83.31 75.68 83.53 84.22 82.43 87.98 80.04 88.80 85.08 82.36 93.60 95.56 88.97
Temporal 94.32 50.97 89.83 84.07 92.07 93.10 61.29 96.40 86.78 87.61 96.45 96.91 91.22
Textual Entailment 88.42 72.82 82.27 83.87 87.41 88.59 60.14 97.07 85.20 85.01 96.23 98.43 89.48
@@
Table 6: Mean BERTScore recall values for various types of reasoning types across different models (Column names are abbreviated and abbreviations can be found in Table 4, BOLD values represent maximum value in pre-trained and instruction-tuned models groups respectively).
@@
Domains Pre-trained Models Instruction-tuned Models SOTA Models
G-2B M-7B G-7B L-3-8B F-2-11B G-2B-I P-3-128k-I M-7B-I G-7B-I L-3-8B-I GPT-3.5-T GPT-4o DS-2
Anthropology 92.87 79.08 91.88 93.18 92.51 95.36 44.46 97.22 92.75 93.21 98.68 99.21 97.23
Books 92.14 63.11 86.18 89.65 87.28 94.19 71.47 96.14 86.23 85.14 95.98 96.33 91.09
Captions 92.01 74.75 87.52 86.04 86.05 93.16 68.74 94.11 85.93 84.90 97.04 98.58 89.54
Code 85.24 93.20 86.76 88.73 84.31 98.87 72.86 99.10 85.26 83.26 99.82 99.87 89.79
Commonsense 85.00 52.62 80.50 81.26 84.29 88.77 56.65 94.03 85.54 83.60 96.13 97.82 91.51
Computer Science 84.76 71.51 82.50 80.65 84.83 87.58 84.95 88.92 85.56 84.23 88.68 89.30 89.12
Debatepedia 91.13 80.81 76.56 87.98 87.75 86.02 46.60 99.18 84.16 85.73 99.18 99.99 87.49
Dialogue 89.15 62.18 84.72 86.33 88.17 89.21 69.98 96.53 84.47 84.24 96.31 98.00 90.56
Economics 82.17 81.28 82.92 80.23 94.22 87.85 85.55 97.90 84.88 86.14 99.93 99.91 99.41
English Exams 93.52 90.02 86.47 89.76 87.45 93.72 73.31 95.29 89.25 87.84 95.69 95.79 94.51
Fiction 93.22 85.01 90.11 93.25 93.39 93.53 70.48 99.13 88.06 87.86 98.15 99.66 92.54
Food 84.96 80.87 91.03 89.24 84.15 86.07 80.49 93.16 84.44 84.82 94.96 95.94 91.87
Formal logic 82.49 82.97 83.44 89.21 83.58 93.98 80.21 99.84 84.36 84.64 87.72 99.86 85.37
Government and Politics 90.65 74.02 86.48 91.09 92.27 90.52 83.96 94.87 86.56 85.86 94.55 96.06 90.30
History 93.89 81.19 90.68 93.66 94.60 93.36 58.20 98.58 89.66 89.89 98.06 99.51 93.08
Justice 92.87 79.08 91.88 93.18 92.51 95.36 44.46 97.22 92.75 93.21 98.68 99.21 97.23
Knowledge Base 89.95 79.40 84.38 85.13 87.31 90.15 88.83 92.69 87.73 86.26 91.96 92.10 92.23
Law 94.23 77.61 89.15 92.57 94.85 93.24 70.70 97.77 88.26 88.61 96.52 98.20 91.49
Linguistics 85.24 93.20 86.76 88.73 84.31 98.87 72.86 99.10 85.26 83.26 99.82 99.87 89.79
Miscellaneous 88.27 64.39 85.86 82.09 85.67 87.06 47.77 96.43 85.44 84.83 96.39 97.52 91.03
Movies 97.80 85.15 94.69 91.53 91.76 95.14 86.13 95.86 83.95 87.34 99.96 99.97 95.89
Narrative 85.40 64.46 81.93 75.63 85.82 87.21 72.76 91.35 83.81 84.12 92.42 93.74 89.33
Natural Science 84.77 77.56 81.85 86.30 91.05 89.25 50.39 95.15 86.39 84.98 94.47 96.41 88.16
News 88.13 70.55 79.73 86.21 87.47 89.14 58.97 90.97 87.35 86.59 90.95 91.36 90.89
Nutrition 86.16 85.18 94.79 92.12 82.84 90.11 84.73 96.79 84.39 86.16 99.50 99.98 93.33
Professions 92.14 95.71 87.32 83.65 82.81 94.68 79.38 94.87 85.49 82.86 95.86 99.25 96.74
Public Places 88.21 61.10 83.44 87.58 87.28 88.96 77.17 90.07 86.43 85.63 93.30 94.92 90.18
Reviews 86.18 74.69 79.74 85.11 87.55 87.51 80.11 89.07 86.13 85.34 89.00 89.59 89.51
School Science Textbooks 92.87 79.08 91.88 93.18 92.51 95.36 44.46 97.22 92.75 93.21 98.68 99.21 97.23
Scientific Research Papers 84.26 76.68 81.28 80.06 85.72 86.89 73.65 86.84 85.22 83.94 89.48 89.88 89.77
Social Media 85.52 84.08 78.21 89.79 91.51 90.87 56.95 96.03 85.43 86.06 95.57 96.09 89.48
Sports 78.82 88.07 80.50 82.61 82.65 80.01 86.04 89.41 81.89 84.15 87.97 88.52 88.67
Statistics 78.82 88.07 80.50 82.61 82.65 80.01 86.04 89.41 81.89 84.15 87.97 88.52 88.67
Story 89.49 71.65 82.21 67.79 89.23 88.29 59.82 92.66 85.25 85.18 91.83 93.81 89.98
Web 87.18 69.98 83.53 82.49 90.95 90.17 70.92 94.90 86.65 86.78 96.85 97.31 91.62
Wikipedia 86.74 55.59 80.90 84.41 88.66 89.03 68.88 95.12 85.72 85.78 95.71 96.03 92.27
@@
Appendix C Qualitative Analyses of Generated Outputs
C.1 Qualitative Examples for Mistral-7B-I
Task Type Domain R. Type Instance Candidate Reference
C.2 Explaining Incorrect Responses of Pre-trained models
In Section 3.7 and Appendix B.1, we observed that even the best pre-trained models are not able to match the performance of IT models on SOTA models. While there was a theoretical reasoning, in this appendix, we will take some qualitative observation of outputs and focus on explainability of errors for the two best performing pre-trained models - Gemma-2B and Falcon-2-11B using the best prompt style.
Appendix D Comparison of Decoding Techniques
Model Name Greedy top-k top-p
$k=10$ $p=0.9$
Gemma-2B 86.41 86.16 86.38
Mistral-7B-v0.3 50.42 50.42 50.42
Gemma-7B 81.06 81.02 81.02
Meta-Llama-3-8B 82.73 83.12 83.05
Falcon-2-11B 86.18 86.12 86.12
Gemma-2B-I 87.96 88.01 87.91
Phi-3-mini-128k-I 55.39 55.43 55.43
Mistral-7B-I-v0.3 88.29 88.29 88.29
Gemma-7B-I 85.14 85.15 85.15
Meta-Llama-3-8B-I 84.11 84.12 84.18
Appendix E Prompt Analyses

### v2

@@ Appendix B Aspect-level Analyses
In this appendix, we report results of all 14 LMs (5 pre-trained, 5 IT and 4 SOTA models that we compared our work to) on all entities of all three aspects present in the test set of the dataset. It includes the ones not covered in Section 2.3, but were available in the test-set of Super-Natural Instructions (Wang et al., 2022), with English as the input and output languages. Note, we also provide the results SOTA models for comparisons. Table 4 reports the results for all task types, Table 6 reports the results on all application domains and Table 5 for all reasoning types. Note, we abbreviate the model names at some places in the columns of these tables. The abbreviations and full model names can be found in Table 3.
In all our analyses, each domain has been considered independent, which is not always the case. There can be some tasks which can be classified into two aspects, like title generation for News articles will belong to title generation task type and News domain. However, in the dataset, there are many such pairwise aspects that do not contain any tasks, and for most of the ones that were present, Mistral-7B-I was the best model. Thus, we are not reporting the tabulated results for aspects considered pairwise considering the sparsity and repetitiveness of such a dense table. This is also discussed in Section 6.
@@ G-7B-I Gemma-7B-I
Task Type # Inst. Pre-trained Models Instruction-tuned Models SOTA Models
S-1.7B G-2B G-7B L-3-8B F-2-11B G-2B-I S-1.7B-I M-7B-I G-7B-I L-3-8B-I GPT-4o-m GPT-4o DS-2 Gem-1.5-Pro
Answerability Classification 1300 84.42 88.80 88.76 87.50 88.29 91.87 90.70 97.39 86.21 85.98 97.82 98.87 91.78 90.93
Cause Effect Classification 700 84.09 86.65 82.16 82.30 88.00 88.03 87.79 94.40 85.56 85.73 95.96 96.59 91.58 90.12
Coreference Resolution 1400 83.03 85.26 75.20 78.76 84.70 88.00 84.68 95.85 84.08 83.27 95.88 97.99 93.12 90.25
Data to Text 826 85.25 87.19 82.01 86.71 88.53 87.72 88.30 90.83 86.14 85.87 91.55 89.62 91.73 89.22
Dialogue Act Recognition 700 82.87 86.16 84.61 83.70 86.60 88.04 88.42 97.04 84.28 83.66 97.74 97.70 91.00 91.16
Grammar Error Correction 100 90.75 93.52 86.47 89.76 87.45 93.72 94.29 95.29 89.25 87.84 95.67 95.79 94.51 94.13
Keyword Tagging 500 82.97 84.23 75.17 83.97 86.03 86.01 85.75 88.52 84.42 83.64 89.13 91.43 89.57 86.01
Overlap Extraction 200 84.91 85.03 84.03 85.27 86.66 86.34 85.82 90.69 85.40 83.83 90.18 92.16 90.04 88.18
Question Rewriting 1100 87.01 90.05 87.81 87.21 89.41 91.05 92.43 93.17 87.80 86.85 93.58 94.37 93.46 92.11
Textual Entailment 2400 82.83 87.91 82.52 83.96 86.61 88.43 86.93 97.07 84.76 84.41 97.00 99.04 89.44 88.15
Title Generation 1784 83.97 84.81 75.57 77.12 86.61 86.67 84.88 89.43 85.44 84.45 88.35 90.01 88.95 87.23
Word Analogy 800 83.93 82.67 83.35 83.22 81.15 88.35 83.63 89.00 85.17 81.53 93.37 97.37 88.44 89.37
@@
Task Type # Inst. Pre-trained Models Instruction-tuned Models SOTA Models
S-1.7B G-2B G-7B L-3-8B F-2-11B S-1.7B-I G-2B-I M-7B-I G-7B-I L-3-8B-I GPT-4o-m GPT-4o DS-2 Gem-1.5-Pro
Abductive 200 85.91 87.58 86.86 86.08 89.67 90.51 88.40 91.36 87.54 87.43 91.12 92.11 91.08 90.71
Analogical 900 83.92 83.02 83.71 83.76 81.73 84.05 88.65 89.15 85.32 81.92 93.07 96.95 88.91 89.58
Causal 800 83.62 87.00 81.19 79.85 86.90 85.21 85.18 93.35 84.78 84.65 95.58 95.88 90.16 89.35
Commonsense 3000 83.80 86.91 83.75 82.26 85.61 87.31 88.96 94.31 85.07 83.89 95.65 97.82 91.35 89.60
Cross-document 200 83.82 94.98 88.45 80.34 92.52 92.88 92.94 95.33 84.15 84.73 96.24 95.53 89.10 88.06
Deductive 200 82.97 91.14 78.63 85.61 88.22 84.58 89.78 92.70 83.73 84.49 91.94 94.75 88.18 88.60
Discrete 200 83.82 94.98 88.45 80.34 92.52 92.88 92.94 95.33 84.15 84.73 96.23 95.53 89.10 88.06
Logical 100 80.87 83.58 77.46 80.16 79.78 82.16 86.13 96.19 81.19 80.50 97.26 99.12 88.09 87.42
Multihop 226 81.65 94.98 86.04 79.84 92.07 88.62 92.94 94.11 83.22 83.82 92.52 94.08 89.01 86.26
Numerical 200 83.82 94.98 88.45 80.34 92.52 92.88 92.94 95.33 84.15 84.73 96.24 95.53 89.10 88.06
Quantitative 300 83.38 88.67 83.60 78.10 86.59 84.11 84.59 93.96 84.80 84.46 94.26 95.19 89.15 87.64
Reasoning on Actions 300 84.96 85.03 82.07 86.80 89.73 85.94 87.18 92.31 86.03 85.12 89.95 92.78 92.48 89.37
Reasoning on Objects 100 88.44 90.79 86.00 91.66 92.66 92.44 91.81 93.08 87.16 87.68 93.95 93.64 93.12 93.25
Social Interactions 700 82.44 84.63 83.36 86.33 86.14 86.44 87.61 93.91 84.52 84.33 95.57 96.55 92.32 91.24
Relational 1100 83.79 83.31 83.53 84.22 82.43 83.94 87.98 88.80 85.08 82.36 92.40 95.56 88.97 89.57
Temporal 300 86.34 94.32 89.83 84.07 92.07 93.98 93.10 96.40 86.78 87.61 96.36 96.91 91.22 90.54
Textual Entailment 2400 83.35 88.42 82.27 83.87 87.41 87.45 88.59 97.07 85.20 85.01 96.64 98.43 89.48 88.71
@@
Table 5: Mean BERTScore recall values for various reasoning types across different models (Column names are abbreviated and abbreviations can be found in Table 3, BOLD values represent the maximum value in pre-trained and instruction-tuned models groups respectively).
@@
Task Type # Inst. Pre-trained Models Instruction-tuned Models SOTA Models
S-1.7B G-2B G-7B L-3-8B F-2-11B S-1.7B-I G-2B-I M-7B-I G-7B-I L-3-8B-I GPT-4o-m GPT-4o DS-2 Gem-1.5-Pro
Anthropology 200 91.68 92.87 91.88 93.18 92.51 95.67 95.36 97.22 92.75 93.21 96.62 99.21 97.23 95.73
Books 300 84.92 92.14 86.18 89.65 87.28 91.42 94.19 96.14 86.23 85.14 96.30 96.33 91.09 89.12
Captions 700 83.94 92.01 87.52 86.04 86.05 90.18 93.16 94.11 85.93 84.90 98.08 98.58 89.54 89.62
Code 100 82.36 85.24 86.76 88.73 84.31 86.23 98.87 99.10 85.26 83.26 99.84 99.87 89.79 89.39
Commonsense 2500 84.33 85.00 80.50 81.26 84.29 85.82 88.77 94.03 85.54 83.60 95.86 97.82 91.51 90.84
Computer Science 100 81.80 84.76 82.50 80.65 84.83 84.21 87.58 88.92 85.56 84.23 86.99 89.30 89.12 85.16
Debatepedia 100 84.44 91.13 76.56 87.98 87.75 83.90 86.02 99.18 84.16 85.73 92.95 99.99 87.49 86.06
Dialogue 1900 82.97 89.15 84.72 86.33 88.17 88.49 89.21 96.53 84.47 84.24 96.33 98.00 90.56 89.25
Economics 100 83.31 82.17 82.92 80.23 94.22 84.46 87.85 97.90 84.88 86.14 84.54 99.91 99.41 88.57
English Exams 100 90.75 93.52 86.47 89.76 87.45 94.29 93.72 95.29 89.25 87.84 95.67 95.79 94.51 94.13
Fiction 700 86.53 93.22 90.11 93.25 93.39 94.50 93.53 99.13 88.06 87.86 97.15 99.66 92.54 89.63
Food 200 82.67 84.96 91.03 89.24 84.15 86.30 86.07 93.16 84.44 84.82 94.84 95.94 91.87 89.78
Formal logic 100 83.43 82.49 83.44 89.21 83.58 83.34 93.98 99.84 84.36 84.64 99.89 99.86 85.37 89.79
Government and Politics 800 84.78 90.65 86.48 91.09 92.27 90.46 90.52 94.87 86.56 85.86 93.74 96.06 90.30 88.34
History 800 87.98 93.89 90.68 93.66 94.60 95.58 93.36 98.58 89.66 89.89 96.67 99.51 93.08 91.22
Justice 200 91.68 92.87 91.88 93.18 92.51 95.67 95.36 97.22 92.75 93.21 96.63 99.21 97.23 95.73
Knowledge Base 100 85.91 89.95 84.38 85.13 87.31 91.83 90.15 92.69 87.73 86.26 91.96 92.10 92.23 92.17
Law 700 86.59 94.23 89.15 92.57 94.85 93.79 93.24 97.77 88.26 95.29 96.52 98.20 91.49
Linguistics 100 82.36 85.24 86.76 88.73 84.31 86.23 98.87 99.10 85.26 83.26 99.84 99.87 89.79 89.39
Miscellaneous 800 84.23 88.27 85.86 82.09 85.67 85.66 87.06 96.43 85.44 84.83 96.64 97.52 91.03 89.71
Movies 100 84.71 97.80 94.69 91.53 91.76 98.77 95.14 95.86 83.95 87.34 99.97 99.97 95.89 90.49
Narrative 800 82.90 85.40 81.93 75.63 85.82 85.98 87.21 91.35 83.81 84.12 91.21 93.74 89.33 88.12
Natural Science 400 84.44 84.77 81.85 86.30 91.05 85.84 89.25 95.15 86.39 84.98 94.86 96.41 88.16 89.92
News 726 85.69 88.13 79.73 86.21 87.47 86.99 89.14 90.97 87.35 86.59 89.80 91.36 90.89 89.06
Nutrition 100 83.24 86.16 94.79 92.12 82.84 88.81 90.11 96.79 84.39 86.16 99.97 99.98 93.33 90.50
Professions 100 83.19 92.14 87.32 83.65 82.81 87.89 94.68 94.87 85.49 82.86 91.16 99.25 96.74 90.35
Public Places 300 86.32 88.21 83.44 87.58 87.28 88.08 88.96 90.07 86.43 85.63 91.35 94.92 90.18 90.94
Reviews 300 85.04 86.18 79.74 85.11 87.55 87.39 87.51 89.07 86.13 85.34 88.19 89.59 89.51 87.78
School Science Textbooks 200 91.68 92.87 91.88 93.18 92.51 95.67 95.36 97.22 92.75 93.21 96.63 99.21 97.23 95.73
Scientific Research Papers 400 82.41 84.26 81.28 80.06 85.72 86.13 86.89 86.84 85.22 83.94 89.29 89.88 89.77 85.34
Social Media 200 84.04 85.52 78.21 89.79 91.51 86.19 90.87 96.03 85.43 86.06 95.75 96.09 89.48 90.06
Sports 26 79.10 78.82 80.50 82.61 82.65 77.75 80.01 89.41 81.89 84.15 85.08 88.52 88.67 82.65
Statistics 26 79.10 78.82 80.50 82.61 82.65 77.75 80.01 89.41 81.89 84.15 85.08 88.52 88.67 82.65
Story 500 83.73 89.49 82.21 67.79 89.23 84.78 88.29 92.66 85.25 85.18 90.67 93.81 89.98 88.25
Web 400 86.67 87.18 83.53 82.49 90.95 89.79 90.17 94.90 86.65 86.78 95.22 97.31 91.62 90.46
Wikipedia 2184 84.60 86.74 80.90 84.41 88.66 88.75 89.03 95.12 85.72 85.78 94.46 96.03 92.27 90.45
@@
Appendix C LM-level Results
Model # Params Def Ex. R-1 R-2 R-L METEOR BERTScore P/R/F1
SmolLM-1.7B 1.71B 2 2.50 1.07 2.27 4.92 67.34/83.71/74.56
Gemma-2B 2.51B 4 22.04 7.88 21.23 18.12 78.22/86.41/81.88
Gemma-2-2B 2.61B 0 7.56 2.18 7.21 9.43 70.29/83.66/76.23
Mistral-7B 7.25B 8 1.17 0.54 1.08 1.99 49.25/58.41/53.40
Gemma-7B 8.54B 0 18.17 5.89 17.49 16.14 71.86/81.06/75.94
Llama-3-8B 8.03B 0 16.38 5.35 15.30 14.96 75.52/82.73/78.80
Falcon-2-11B 11.1B 8 16.88 6.46 16.01 16.45 79.65/86.18/82.72
SmolLM-1.7B-I 1.71B 2 20.22 7.59 19.03 18.78 80.34/86.66/83.24
Gemma-2B-I 2.51B 2 27.56 8.08 26.24 20.62 84.56/88.06/86.19
Gemma-2-2B-I 2.61B × 0 3.45 1.35 2.99 5.60 73.91/82.76/78.05
Mistral-7B-I 7.25B 8 51.96 14.67 50.12 35.55 91.29/93.76/92.39
Gemma-7B-I 8.54B 0 8.64 3.23 7.96 12.57 78.18/85.14/81.48
Llama-3-8B-I 8.03B × 8 4.68 2.19 4.23 8.31 74.23/84.33/78.89
In Table 7, we report the best prompt style at the LM-level, abstracting all analyses at aspect-level with different performance metrics like ROUGE 1/2/L (Lin, 2004), METEOR (Banerjee and Lavie, 2005) and BERTScore P/R/F1 (Zhang et al., 2019) for reference.
Here we also include the results on Gemma-2-2B, Gemma-2-2B-I (Team et al., 2024c) and Mistral-7B-v0.3 (Jiang et al., 2023). From the results, it is visible why we ignored these models from main analysis. The Gemma-2 family is performing less compared to Gemma, and we wanted to keep a wide LM family for analysis. For Mistral, it was just underperforming. So, for brevity, we excluded them.
Appendix D Prompt Analyses